OpenStudy (anonymous):

Write the slope-intercept equation of the line perpendicular to y = 5/2 x - 2, which passes through the point (0, 2). A. y = 5x + 2 B. y = -2/5x + 2 C. y = 2x + 5 D. y = 1/2x + 5

OpenStudy (anonymous):

hmmm lets see what we have here

OpenStudy (aroub):

y=mx+b , m=-2/5 , (0,2) 2=-2/5(0)+b 2=b => y=-2/5x+2